### v3

# Clutter-Aware Target Detection for ISAC in a Millimeter-Wave Cell-Free Massive MIMO System

Abstract
In this paper, we investigate the performance of an integrated sensing and communication (ISAC) system within a cell-free massive multiple-input multiple-output (MIMO) system. Each access point (AP) operates in the millimeter-wave (mmWave) frequency band. The APs jointly serve the user equipments (UEs) in the downlink while simultaneously detecting a target through dedicated sensing beams directed toward a reconfigurable intelligent surface (RIS). Although the AP-RIS, RIS-target, and AP-target channels have both line-of-sight (LoS) and non-line-of-sight (NLoS) parts, only knowledge of the LoS paths is assumed to be available. A key contribution of this study is the consideration of clutter, which degrades the target detection performance if not handled. We propose an algorithm to alternatively optimize the transmit power allocation and the RIS phase-shift matrix, maximizing the target signal-to-clutter-plus-noise ratio (SCNR) while ensuring a minimum signal-to-interference-plus-noise ratio (SINR) for the UEs. Numerical results demonstrate that exploiting clutter subspace significantly enhances detection probability, particularly at high clutter-to-noise ratios, and reveal that an increased number of transmit side clusters impairs detection performance. Finally, we highlight the performance gains achieved using a dedicated sensing stream.
Index Terms:
RIS, ISAC, cell-free massive MIMO, mmWave.
I Introduction
Cell-free massive multiple-input multiple-output (MIMO) is a user-centric network infrastructure where a set of distributed access points serve multiple user equipments on the same time-frequency resources using joint processing techniques [1]. Over nearly a decade of research, it has been shown that cell-free massive MIMO improves both spectral and energy efficiency by providing macro-diversity and enhanced interference management compared to traditional cellular setups. As a result, it has become one of the key technology components considered for sixth-generation (6G) wireless networks [2]. Recently, cell-free massive MIMO architecture has also demonstrated potential benefits in integrated sensing and communications (ISAC) [3, 4]. ISAC allows for the efficient use of hardware and spectral resources through integration and coordination gains, compared to separate communication and sensing architectures [5]. Cell-free massive MIMO facilitates multi-static sensing through a central processing unit (CPU), to which the APs are connected via fronthaul links. Beyond cell-free massive MIMO and ISAC, another key 6G technology is reconfigurable intelligent surface (RIS)-aided communications. RISs have emerged as a promising technology in ISAC applications, enhancing the available spatial degrees of freedom. In [6], the sum rate of communication UEs is maximized under a worst-case sensing signal-to-noise ratio (SNR) constraint. An often overlooked aspect in ISAC research is the effect of clutter on target detection, which can severely degrade system performance if not properly addressed. Recent studies, such as [7], illustrate the impact of clutter in monostatic ISAC settings, underscoring the significance of clutter awareness at the receiving APs for reliable target detection.
Refer to captionRISCloud unitAP $r\in\mathcal{R}$TargetClutterAP $t\in\mathcal{T}$UE $k\in\mathcal{K}$$\mathbf{g}_{k,t}$$\mathbf{H}_{t}$$\mathbf{c}$$\mathbf{b}_{r}$ $\overline{\mathbf{H}}_{t}$$\overline{\mathbf{c}}$$\overline{\mathbf{b}}_{r}$ cluster
Figure 1: Depiction of a cell-free massive MIMO ISAC network performing multi-static sensing.
The contributions of this paper are outlined as follows
  • we investigate the ability of a RIS to improve target detection under transmitter-target obstruction. The need for precise channel state information (CSI) for target detection and the requirement for sensing-specific streams are explored within a millimeter-wave (mmWave) channel architecture.
  • A signal-to-clutter-and-noise ratio (SCNR)-maximizing alternating optimization (AO) algorithm is proposed, where the system assumes a line-of-sight (LoS) structure for the sensing links and optimizes the transmit power allocation and the RIS phase-shift matrix
  • numerical simulations reveal that clutter awareness and dedicated sensing streams significantly enhance the detection probability.
II System Model
We consider the RIS-assisted cell-free massive MIMO ISAC network shown in Fig. 1. A set $\mathcal{T}$ of $T$ APs, each equipped with $M$ transmit antennas, serves $K$ single-antenna UEs in the downlink while sensing the potential presence of a target at a known position. Sensing is carried out over $\tau$ symbols belonging to the same channel coherence block. A second set $\mathcal{R}$ of $R$ APs collects the target echoes and performs target detection. The choice of geographically separated transmitting and receiving APs is motivated by the implementation challenges of full duplex operation, especially in terms of self-interference [8], and the interest in this setup for cell-free massive MIMO communications. We assume that the direct channel between the APs in $\mathcal{T}$ and the target is blocked and, thus, the formers reach the target through a RIS equipped with $N$ reflective elements, whose phase shifts are described by the vector $\boldsymbol{\theta}=[\theta_{1},\dots,\theta_{N}]^{\top}$, where $|\theta_{n}|=1$ for $n=1,\dots,N$. Each AP $t\in\mathcal{T}$ is connected to the RIS via the channel $\mathbf{H}_{t}\in\mathbb{C}^{N\times M}$, the reflected path connecting the RIS to the target is denoted by $\mathbf{c}\in\mathbb{C}^{N}$ and the UEs are connected to the APs in $\mathcal{T}$ via the channel $\{\mathbf{g}_{k,t}\}_{\forall k,t}\in\mathbb{C}^{M}$. The receiving APs in $\mathcal{R}$ are connected to the target via the channels $\{\mathbf{b}_{r}\}_{\forall r\in\mathcal{R}}\in\mathbb{C}^{M}$.
II-A Channel models
Each of the previously mentioned channels is modelled in accordance with [9], which accurately captures the spatial sparsity that characterizes mmWave channels. Each channel consists of contributions from $C_{i}$ scattering clusters, i.e.,
$\mathbf{b}_{r}=\underbrace{\sqrt{\beta^{2}_{0,r}}\hskip 2.84526pt\mathbf{a}_{M}(\boldsymbol{\omega}_{0,r}^{\textrm{RX}})}_{\overline{\mathbf{b}}_{r}}+\sqrt{\frac{1}{C_{1}}}\sum_{n=1}^{C_{1}}\alpha_{n,r}\hskip 1.42262pt\mathbf{a}_{M}(\boldsymbol{\omega}_{n,r}^{\textrm{RX}}),$ (1)
$\mathbf{c}=\underbrace{\sqrt{\beta_{0}^{2}}\hskip 2.84526pt\mathbf{a}_{N}(\boldsymbol{\omega}_{0}^{\textrm{Tg}})}_{\overline{\mathbf{c}}}+\sqrt{\frac{1}{C_{2}}}\sum_{n=1}^{C_{2}}\alpha_{n}\hskip 1.42262pt\mathbf{a}_{N}(\boldsymbol{\omega}_{n}^{\textrm{Tg}}),$ (2)
$\mathbf{g}_{k,t}=\sqrt{\frac{1}{C_{3}}}\sum_{n=1}^{C_{3}}\alpha_{n,k,t}\hskip 1.42262pt\mathbf{a}_{M}(\boldsymbol{\omega}_{n,k,t}^{\textrm{UE}}),$ (3)
$\mathbf{H}_{t}=\overbrace{\sqrt{\beta_{0,t}^{2}}\hskip 2.84526pt\mathbf{a}_{N}(\boldsymbol{\omega}_{0,t}^{\textrm{RIS}})\mathbf{a}_{M}^{\top}(\boldsymbol{\omega}_{0,t}^{\textrm{TX}})}^{\overline{\mathbf{H}}_{t}}$
[…]
where the overline superscript indicates the LoS paths and the other terms correspond to the NLoS parts. The coefficients $\beta_{0,r}^{2}$, $\beta_{0}^{2}$, and $\beta_{0,t}^{2}$ are the corresponding LoS channel gains. We now remove the additional subscripts and superscripts except for $n$ as the following consideration applies regardless of the additional subscripts. Here, $\alpha_{n}$ is the complex channel gain of the $n$-th cluster. We assume each cluster scatters a sufficient number of rays such that $\alpha_{n}\sim\mathcal{CN}(0,\beta_{n}^{2})$ and the fading is mutually independent. The coefficients $\beta_{n}^{2}$ represent the path loss associated with the respective clusters. The geometry of the $n$-th cluster is described by $\boldsymbol{\omega}_{n}=[\psi_{n},\phi_{n}]$, where $\psi_{n}$ is the azimuth angle and $\phi_{n}$ is the elevation angle. Each AP is equipped with a uniform planar square array with a half-wavelength vertical and horizontal inter-element spacing such that the entries of $M$-dimensional steering vector are defined as $[\mathbf{a}_{M}(\boldsymbol{\omega}_{n})]_{m_{h}m_{v}}=e^{-j\pi(m_{h}\sin(\psi_{n})\cos(\phi_{n})+m_{v}\sin(\phi_{n}))}$,for $m_{h},m_{v}=0,\dots,\sqrt{M}-1$. We assume that perfect CSI regarding $\{\mathbf{g}_{k,t}\}_{\forall k,t}$ is available at the APs, as this is easily obtainable through standard estimation techniques. As for the sensing channels, the transmit APs know the position of the RIS and the target and, thus, the corresponding LoS channels.
II-B Communication observation model
We assume that the APs are equipped with fully digital beamforming hardware: $\mathbf{f}_{l,t}$ is the precoding vector for stream $l$ and AP $t$, and $\mathbf{F}_{t}=[\mathbf{f}_{1,t},\ldots,\mathbf{f}_{L,t}]\in\mathbb{C}^{M\times L}$, where $L=K+1$. Here, we allocate one beam to each UE and an additional one, denoted by the index $L$, for sensing purposes, bringing the total number of digital beams to $L=K+1$. This choice is motivated by the fact that the UEs and target typically are at very different positions, and catering solely to the communication UEs might drastically reduce the system’s sensing performance.
The received downlink signal at UE $k$ during timeslot $\iota$ is defined as
$y_{k}[\iota]$ $=\sum_{t\in\mathcal{T}}\mathbf{g}_{k,t}^{\mathsf{H}}\mathbf{s}_{t}[\iota]+w_{k}[\iota]=\sum_{t\in\mathcal{T}}\Big{(}\underbrace{\sqrt{\rho_{k}}\mathbf{g}_{k,t}^{\mathsf{H}}\mathbf{f}_{k,t}x_{k}[\iota]}_{\text{desired signal}}$ (5)
$+\underbrace{\sum_{k^{\prime}\in\mathcal{K}\setminus\{k\}}\sqrt{\rho_{k^{\prime}}}\mathbf{g}_{k,t}^{\mathsf{H}}\mathbf{f}_{k^{\prime},t}x_{k^{\prime}}[\iota]}_{\text{multi-user interf. (MUI)}}+\underbrace{\sqrt{\rho_{L}}\mathbf{g}_{k,t}^{\mathsf{H}}\mathbf{f}_{L,t}x_{L}[\iota]}_{\text{sensing interf.}}\Big{)}+w_{k}[\iota],$
with $\mathbf{s}_{t}[\iota]=\mathbf{F}_{t}\mathbf{X}[\iota]\boldsymbol{\rho}$, where the matrix $\mathbf{X}[\iota]=\text{diag}(x_{1}[\iota],\ldots,x_{L}[\iota])$ is a diagonal matrix containing the downlink communication symbols and sensing symbol transmitted during timeslot $\iota$, where $\mathbb{E}\{|x_{k}[\iota]|^{2}\}=1$. The vector $\boldsymbol{\rho}=[\sqrt{\rho_{1}},\dots,\sqrt{\rho_{L}}]^{\top}$ contains the square root of the power allocated to each beam. Finally, $w_{k}[\iota]\sim\mathcal{CN}(0,\sigma_{k}^{2})$. In line with the centralized operation of cell-free massive MIMO, the communication precoding vectors are based on the transmit APs’ CSI [3] while the sensing precoder is based on the assumed LoS structure of the sensing channels and the known target location. To this end, we define $\mathbf{g}_{k}=[\mathbf{g}_{k,1}^{\top},\dots,\mathbf{g}_{k,T}^{\top}]^{\top}\in\mathbb{C}^{MT}$ and $\mathbf{f}_{k}=[\mathbf{f}_{k,1}^{\top},\dots,\mathbf{f}_{k,T}^{\top}]^{\top}\in\mathbb{C}^{MT}$. The communication precoding vectors for each UE are chosen using regularized zero forcing (RZF) as $\mathbf{f}_{k}=\overline{\mathbf{f}}_{k}/\left\lVert\overline{\mathbf{f}}_{k}\right\rVert$, where
$\overline{\mathbf{f}}_{k}=\left(\sum_{i\in\mathcal{K}}\mathbf{g}_{i}\mathbf{g}_{i}^{\mathsf{H}}+\lambda\mathbf{I}_{MT}\right)^{-1}\mathbf{g}_{k}$ (6)
where $\lambda$ is a regularization parameter. Let us define the matrix $\mathbf{G}=[\mathbf{g}_{1},\dots,\mathbf{g}_{K}]\in\mathbb{C}^{MT\times K}$ with all the UE channels. Then the sensing stream precoder is defined as $\mathbf{f}_{L,t}=\overline{\mathbf{f}}_{L,t}/\left\lVert\overline{\mathbf{f}}_{L,t}\right\rVert$
$\overline{\mathbf{f}}_{L,t}=\left(\mathbf{I}_{MT}-\mathbf{G}\left(\mathbf{G}^{\mathsf{H}}\mathbf{G}\right)^{\dagger}\mathbf{G}^{\mathsf{H}}\right)\overline{\mathbf{h}}$ (7)
where $\overline{\mathbf{h}}=[\overline{\mathbf{h}}_{1}^{\top},\dots,\overline{\mathbf{h}}_{T}^{\top}]^{\top}$ and $\overline{\mathbf{h}}_{t}=\overline{\mathbf{H}}_{t}^{\mathsf{H}}\boldsymbol{\Theta}\overline{\mathbf{c}}$ with $\boldsymbol{\Theta}=\text{diag}(\boldsymbol{\theta})$. We have projected $\overline{\mathbf{h}}$ onto the null space of the communication UEs’ channels to nullify the interference caused to the UEs. The communication performance is represented by the UEs’ spectral efficiency (SE), defined as ${\rm{SE}}_{k}=\log_{2}(1+{\rm{SINR}}_{k})$: under the assumption of perfect CSI on the UEs channels, the signal-to-interference-plus-noise ratio (SINR) is defined as
${\rm{SINR}}_{k}=\frac{|{\mathbf{g}}_{k}^{\mathsf{H}}\mathbf{f}_{k}\sqrt{\rho_{k}}|^{2}}{\sum_{k^{\prime}\in\mathcal{K}/k}|{\mathbf{g}}_{k}^{\mathsf{H}}\mathbf{f}_{k^{\prime}}\sqrt{\rho_{k^{\prime}}}|^{2}+|{\mathbf{g}}_{k}^{\mathsf{H}}\mathbf{f}_{L}\sqrt{\rho_{L}}|^{2}+\sigma_{k}^{2}}.$ (8)
II-C Sensing observation model
The two-way sensing channel between transmit AP $t$ and receive AP $r$ is defined as $\mathbf{E}_{t,r}=\mathbf{b}_{r}\mathbf{h}_{t}^{\mathsf{H}}\triangleq\mathbf{b}_{r}\left(\mathbf{H}_{t}^{\mathsf{H}}\boldsymbol{\Theta}\mathbf{c}\right)^{\mathsf{H}}.$ The sensing observation of AP $r$ at timeslot $\iota$ is
$\mathbf{y}_{r}[\iota]=\sum_{t\in\mathcal{T}}\xi_{t,r}\mathbf{E}_{t,r}\mathbf{s}_{t}[\iota]+\mathbf{z}_{r}[\iota]+\mathbf{n}_{r}[\iota],\vspace{-2mm}$ (9)
where $\xi_{t,r}$ is the radar cross section (RCS) associated with the sensing path between transmit AP $t$ and receive AP $r$. We adopt the Swerling-I model, meaning that the RCS assumes only one value during the transmission and $\xi_{t,r}\sim\mathcal{CN}(0,\delta_{t,r}^{2})$. We assume that $\{\delta_{t,r}^{2}\}_{\forall t,r}$ are known and that RCSs belonging to different transmit-receiving pairs are statistically uncorrelated, that is $\mathbb{E}\left\{\xi_{a,r}\xi_{a^{\prime},r^{\prime}}^{*}\right\}\neq 0\Leftrightarrow a=a^{\prime},\leavevmode\nobreak\ r=r^{\prime}$. In (9), $\mathbf{n}_{r}[\iota]$ is the receiver noise of AP $r$, assumed to have independent $\mathcal{CN}(0,\sigma^{2})$ entries. The vector $\mathbf{z}_{r}[\iota]$ represents the clutter as observed from AP $r$. The clutter vector is modelled as $\mathbf{z}_{r}[\iota]\sim\mathcal{CN}(\boldsymbol{0},\delta_{\mathbf{z}}^{2}\mathbf{R}_{r})$, where $\delta_{\mathbf{z}}^{2}$ is the power of the clutter cross-section, which also includes the path loss effects, thus defining the clutter-to-noise ratio (CNR) as $\textrm{CNR}=\delta_{\mathbf{z}}^{2}/\sigma^{2}$. The normalized clutter covariance matrix, with a trace equal to $M$, is denoted by $\mathbf{R}_{r}$. This matrix is unknown, but its eigenspace is assumed to be known later in the detector design. For the sake of mathematical tractability, we assume that the clutter observations belonging to different APs and/or different timeslots are statistically uncorrelated. Under the Los assumption for the channels involved in sensing, the target’s SCNR is defined as
$\overline{\textrm{SCNR}}=\frac{\mathbb{\sum}_{\iota=1}^{\tau}\mathbb{\sum}_{r\in\mathcal{R}}\sum_{a\in\mathcal{T}}\delta_{a,r}^{2}\left\lVert\overline{\mathbf{E}}_{a,r}\mathbf{s}_{a}[\iota]\right\rVert^{2}}{R\tau M\left(\sigma^{2}+\delta_{\mathbf{z}}^{2}\right)},$ (10)
where $\overline{\mathbf{E}}_{a,r}=\overline{\mathbf{b}}_{r}\left(\overline{\mathbf{H}}_{a}^{\mathsf{H}}\boldsymbol{\Theta}\overline{\mathbf{c}}\right)^{\mathsf{H}}$.
III Sensing performance optimization
The system aims at jointly maximizing the sensing SCNR (10) while guaranteeing a minimum SINR to all the communication UEs. The optimization problem is stated as follows:
$\underset{\boldsymbol{\theta},\boldsymbol{\rho}}{\mathrm{maximize}}\,\,$ $\leavevmode\nobreak\ \overline{\textrm{SCNR}}$ (11a)
subject to $\textrm{SINR}_{k}\geq\gamma_{k},\quad\forall k\in\mathcal{K},$ (11b)
$\sum_{l=1}^{L}\left\lVert\mathbf{f}_{l,t}\right\rVert^{2}\rho_{l}\leq P_{t},\quad\forall t\in\mathcal{T},$ (11c)
$|\theta_{n}|=1,\quad n=1,\dots,N,$ (11d)
where $\gamma_{k}$ is the SINR threshold of UE $k$ and $P_{t}$ is the available transmit power of the $t$-th AP. This problem is non-convex, mainly due to the coupling between the optimization variables and the non-convex unitary modulus constraints. In this section, we will devise an AO algorithm, which switches between optimizing the power allocation policy and the RIS phase shifts until convergence is achieved.
III-A Power allocation optimization
In this subsection, we introduce the first subproblem of the AO algorithm, where we optimize the power allocation for a fixed RIS configuration. We first reformulate (10) as $\overline{\textrm{SCNR}}=\boldsymbol{\rho}^{\top}\Re(\mathbf{A})\boldsymbol{\rho}$ where
$\mathbf{A}=\frac{\sum_{\iota=1}^{\tau}\mathbb{\sum}_{r\in\mathcal{R}}\sum_{a\in\mathcal{T}}\delta_{a,r}^{2}\mathbf{X}^{\mathsf{H}}[\iota]\mathbf{F}_{a}^{\mathsf{H}}\overline{\mathbf{E}}_{a,r}^{\mathsf{H}}\overline{\mathbf{E}}_{a,r}\mathbf{F}_{a}\mathbf{X}[\iota]}{R\tau M\left(\sigma^{2}+\delta_{\mathbf{z}}^{2}\right)}.$ (12)
The constraints (11b) and (11c) can be reformulated as second order cone (SOC) constraints in terms of $\boldsymbol{\rho}$. The first subproblem of the AO SCNR-maximizing algorithm is defined as
$\underset{\boldsymbol{\rho}}{\mathrm{maximize}}\,\,$ $\leavevmode\nobreak\ \boldsymbol{\rho}^{\top}\Re(\mathbf{A})\boldsymbol{\rho}$ (13a)
subject to $\Big{[}|\mathbf{g}_{k}^{\mathsf{H}}\mathbf{f}_{1}|\sqrt{\rho_{1}}\quad\dots\quad\underbrace{0}_{k}\quad\dots\quad|\mathbf{g}_{k}^{\mathsf{H}}\mathbf{f}_{L}|\sqrt{\rho_{L}}$
$\sigma_{k}\Big{]}\leq(|\mathbf{g}_{k}^{\mathsf{H}}\mathbf{f}_{k}|\sqrt{\rho_{k}})/{\sqrt{\gamma_{k}}},\leavevmode\nobreak\ \forall k\in\mathcal{K}$ (13b)
$\left\lVert\widetilde{\mathbf{F}}_{t}\boldsymbol{\rho}\right\rVert\leq\sqrt{P_{t}},\leavevmode\nobreak\ \forall t\in\mathcal{T}$ (13c)
where $\widetilde{\mathbf{F}}_{t}=\text{diag}\left(\left\lVert\mathbf{f}_{1,t}\right\rVert,\dots,\left\lVert\mathbf{f}_{L,t}\right\rVert\right)$. This problem can be solved using the convex-concave procedure outlined in [3].
III-B RIS phase-shift optimization
We now move on to the second part of the AO, where we optimize $\boldsymbol{\theta}$ when the power variables are fixed. The SCNR can be reformulated as $\overline{\textrm{SCNR}}=\boldsymbol{\theta}^{\mathsf{H}}\mathbf{Q}\boldsymbol{\theta}$, where
$\mathbf{Q}=$ $\frac{1}{R\tau M\left(\sigma^{2}+\delta_{\mathbf{z}}^{2}\right)}\sum_{\iota=1}^{\tau}\sum_{r\in\mathcal{R}}\sum_{a\in\mathcal{T}}\Big{(}\delta_{a,r}^{2}\overline{\mathbf{c}}\overline{\mathbf{b}}_{r}^{\mathsf{H}}\overline{\mathbf{b}}_{r}\overline{\mathbf{c}}^{\mathsf{H}}\Big{)}^{\top}$
$\quad\quad\odot\left(\overline{\mathbf{H}}_{a}\mathbf{s}_{a}[\iota]\mathbf{s}_{a}^{\mathsf{H}}[\iota]\overline{\mathbf{H}}_{a}^{\mathsf{H}}\right).$ (14)
We have used $\text{Tr}(\mathbf{A}\mathbf{X}\mathbf{B}\mathbf{X}^{\mathsf{H}})=\text{vec}^{\mathsf{H}}(\mathbf{X})\left(\mathbf{B}^{\top}\otimes\mathbf{A}\right)\text{vec}(\mathbf{X})$ and $\text{vec}^{\mathsf{H}}(\text{diag}(\mathbf{x}))(\mathbf{A}\otimes\mathbf{B})\text{vec}(\text{diag}(\mathbf{x}))=\mathbf{x}^{\mathsf{H}}(\mathbf{A}\odot\mathbf{B})\mathbf{x}$ from [10]. The RIS phase-shift optimization problem under consideration is still non-convex: We then employ the minorization-maximization (MM) algorithm to maximize a convex lower bound on the objective function [11] around the local point $\boldsymbol{\theta}^{(s)}$. We can now define the second subproblem of the AO optimization as
$\underset{\boldsymbol{\theta},\alpha\geq 0}{\textrm{maximize}}$ $\leavevmode\nobreak\ \alpha$ (15a)
subject to $\alpha\leq 2\Re\left(\boldsymbol{\theta}^{(s)\mathsf{H}}\mathbf{Q}^{(+)}\boldsymbol{\theta}\right)+\boldsymbol{\theta}^{\mathsf{H}}\mathbf{Q}^{(-)}\boldsymbol{\theta}$ (15b)
$\quad\quad-\boldsymbol{\theta}^{(s)\mathsf{H}}\mathbf{Q}^{(+)}\boldsymbol{\theta}^{(s)},$
$|\theta_{n}|\leq 1,\quad n=1,\dots,N,$ (15c)
where $\alpha$ is an auxiliary variable and $\mathbf{Q}^{(+)}$ and $\mathbf{Q}^{(-)}$ represent the positive and negative semidefinite parts of $\mathbf{Q}$. Note that we have relaxed the unit modulus constraint, as it is observed that the solution satisfies this constraint with equality anyway. This is a convex problem and can be solved at each iteration until convergence is achieved. The steps of the AO algorithm are outlined in Algorithm 1.
1:Initialize: Generate $\boldsymbol{\theta}^{(0)}$ randomly, $v\leftarrow 0$
2:repeat
3:   Compute the precoding codebooks with $\boldsymbol{\theta}=\boldsymbol{\theta}^{(v)}$
4:   Obtain $\boldsymbol{\rho}^{(v+1)}$ by solving (13)
5:   Obtain $\boldsymbol{\theta}^{(v+1)}$ by solving (15) iteratively
6:   $v\leftarrow v+1$
7:until Convergence
8:Output: $\boldsymbol{\theta}^{\text{opt}},\boldsymbol{\rho}^{\text{opt}}$
Algorithm 1 SCNR maximizing AO algorithm
IV GLRT Detector
One of the main metrics used to assess the sensing performance is the probability of correct detection, implemented through a generalized likelihood ratio test (GLRT) by extending the detector in [7] to a cell-free massive MIMO system based on the optimization procedure developed in the previous section. We operate under the assumption that the transmit signals are known by the receiving APs, thanks to the cell-free architecture. We define the vector containing the combined sensing observation of all receiving APs during timeslot $\iota$ as
$\mathbf{y}[\iota]=[\mathbf{y}_{1}^{\top}[\iota],\dots,\mathbf{y}_{R}^{\top}[\iota]]^{\top}=\mathbf{V}[\iota]\boldsymbol{\xi}+\mathbf{U}\mathbf{w}[\iota]+\mathbf{n}[\iota],$
$\mathbf{V}[\iota]=\textrm{blkdiag}(\mathbf{V}_{1}[\iota],\dots,\mathbf{V}_{R}[\iota])\in\mathbb{C}^{MR\times TR},$
$\mathbf{V}_{r}[\iota]=\left[\mathbf{E}_{1,r}\mathbf{s}_{1}[\iota],\dots,\mathbf{E}_{T,r}\mathbf{s}_{T}[\iota]\right]\in\mathbb{C}^{M\times T},$
$\mathbf{n}[\iota]=[\mathbf{n}_{1}^{\top}[\iota],\dots,\mathbf{n}_{R}^{\top}[\iota]]^{\top},\leavevmode\nobreak\ \boldsymbol{\xi}=[\xi_{1,1},\dots,\xi_{T,R}]^{\top},$
$\mathbf{U}=\textrm{blkdiag}\left(\mathbf{U}_{1},\dots,\mathbf{U}_{R}\right),\mathbf{w}[\iota]=[\mathbf{w}_{1}^{\top}[\iota],\dots,\mathbf{w}_{R}^{\top}[\iota]]^{\top}.$
The clutter observed by AP $r$ is assumed to exist in a low-rank subspace [12] spanned by the columns of the semi-unitary matrix $\mathbf{U}_{r}\in\mathbb{C}^{M\times i}$, with $i$ being the subspace dimension, obtained from the eigenspace of $\delta_{\mathbf{z}}^{2}\mathbf{R}_{r}$. AP $r$’s clutter is then expressed as $\mathbf{U}_{r}\mathbf{w}_{r}[\iota]$, where the distribution of $\mathbf{w}_{r}[\iota]$ is unknown. The previous observations are generated using the true channels; however, the system does not have perfect CSI on the sensing channels and thus assumes a LoS structure. Therefore, the binary hypotheses being tested are
$\mathcal{H}_{0}:\mathbf{y}[\iota]=\overline{\mathbf{V}}[\iota]\boldsymbol{\xi}+\mathbf{U}\mathbf{w}[\iota]+\mathbf{n}[\iota],\leavevmode\nobreak\ \iota=1,\dots,\tau,$ (16a)
$\mathcal{H}_{1}:\mathbf{y}[\iota]=\mathbf{U}\mathbf{w}[\iota]+\mathbf{n}[\iota],\leavevmode\nobreak\ \iota=1,\dots,\tau,$ (16b)
where $\overline{\mathbf{V}}$ is computed as $\mathbf{V}$ by using $\overline{\mathbf{E}}_{t,r}$. The adoption of a GLRT method is motivated by the unknown distribution of $\mathbf{w}[\iota]$, however, we know that $\mathbf{y}_{\mathcal{H}_{0}}[\iota]-\mathbf{U}\mathbf{w}[\iota]\sim\mathcal{CN}(\mathbf{0},\sigma^{2}\mathbf{I}_{MR})$ and $\mathbf{y}_{\mathcal{H}_{1}}[\iota]-\mathbf{U}\mathbf{w}[\iota]\sim\mathcal{CN}(\mathbf{0},\mathbf{R}[\iota]+\sigma^{2}\mathbf{I}_{MR})$, where $\mathbf{R}[\iota]=\mathbb{E}\left\{\overline{\mathbf{V}}[\iota]\boldsymbol{\xi}\boldsymbol{\xi}^{\mathsf{H}}\overline{\mathbf{V}}^{\mathsf{H}}[\iota]\right\}=\overline{\mathbf{V}}[\iota]\text{diag}(\delta^{2}_{1,1},\dots,\delta^{2}_{T,R})\overline{\mathbf{V}}^{\mathsf{H}}[\iota]$. Following the derivation in [7], the GLRT test can be rewritten as
$\sum_{\iota=1}^{\tau}\mathbf{y}^{\mathsf{H}}[\iota]\Big{(}\boldsymbol{\Xi}[\iota]\mathbf{U}\Big{(}\mathbf{U}^{\mathsf{H}}\boldsymbol{\Xi}[\iota]\mathbf{U}\Big{)}^{-1}\mathbf{U}^{\mathsf{H}}\boldsymbol{\Xi}[\iota]+$
$\sigma^{-2}(\mathbf{I}_{MR}-\mathbf{U}\mathbf{U}^{\mathsf{H}})-\boldsymbol{\Xi}[\iota]\Big{)}\mathbf{y}[\iota]\lessgtr^{\mathcal{H}_{0}}_{\mathcal{H}_{1}}\widetilde{\lambda}_{d}$ (17)
where $\boldsymbol{\Xi}[\iota]=(\mathbf{R}[\iota]+\sigma^{2}\mathbf{I}_{MR})^{-1}$. $\widetilde{\lambda}_{d}=\ln(\lambda_{d})-\ln\left(\det\left(\sigma^{2}\mathbf{I}_{MR}\right)^{\tau}/{\prod_{\iota=1}^{\tau}\det\left(\mathbf{R}[\iota]+\sigma^{2}\mathbf{I}_{MR}\right)}\right)$ is a revised threshold whose value should be chosen according to the desired false alarm probability.
5050505010010010010015015015015000101010102020202030303030x (m)y (m)$\mathcal{T}$$\mathcal{R}$
Figure 2: 2D locations in the simulated scenario.
V Numerical Results
[…]
Figure 3: Probability of detection vs RCS power with CNR = $20$ dB.
[…]
Figure 4: Probability of detectionvs the RCS power with a CNR=$40$ dB.
V-A Clutter awareness and cluster number’s impact
Fig. 3 shows this probability of detection vs. the RCS power, where we assumed that $\delta_{1,1}^{2}=\dots=\delta_{T,R}^{2}=\delta^{2}$ for a CNR of $20$ dB. We can see that a clutter-aware system outperforms a clutter-unaware one: The clutter-unaware detector is defined as $\mathbf{T}[\iota]=\sigma^{-2}\mathbf{I}_{MR}-\boldsymbol{\Xi}[\iota]$. When $C_{2}=C_{3}=C_{4}=20$, denoted by $C_{234}$, the LoS assumption is no longer realistic, thus generating a performance decrease. Our optimized model greatly outperforms the benchmark approach, consisting of a random RIS phase-shift matrix and an equal power allocation between the beams. Fig. 4 shows what happens when the CNR rises to $40$ dB. We see that clutter awareness rises in importance, widening the gap between the clutter-aware and unaware detectors.
[…]
Figure 5: Impact of a dedicated sensing stream onto the probability of detection with a CNR=$20$ dB.
V-B Do we need dedicated sensing symbols?
We have analyzed a system without sensing-dedicated precoding. Fig. 5 compares the detection performance of systems with and without dedicated sensing precoding, both in the presence and absence of clutter awareness. We can immediately see that removing the sensing streams implies a performance degradation. Interestingly, the system without sensing dedicated precoding seems to suffer less from the lack of clutter awareness at a low RCS value.
VI Conclusions
We have investigated a cell-free massive MIMO ISAC network jointly serving multiple communication UEs while detecting a potential target through an RIS in a multi-static fashion. To this end, the transmit power allocation and the RIS phase-shift matrix are optimized to increase the target detection probability. Numerical results show that clutter awareness plays a crucial role in the detection performance and that including a dedicated sensing stream benefits the system. The mmWave channel structure allows the system to assume LoS channels and still achieve good performances when the channel is composed of a high number of clusters concentrated in a small angular range.